Problem 6-5
Expected Return: Discrete Distribution

A stock's return has the following distribution:


Demand for the
Company's Products
Probability of This
Demand Occurring
Rate of Return if This  
Demand Occurs (%)
Weak 0.1 -50%
Below average 0.2 -6
Average 0.4 9
Above average 0.2 30
Strong 0.1 75
1.0

Calculate the stock's expected return. Round your answer to two decimal places.
%?

Calculate the standard deviation. Round your answer to two decimal places?
Answer #1

Expected return=Respective return*Respective probability

=(0.1*-50)+(0.2*-6)+(0.4*9)+(0.2*30)+(0.1*75)

=10.9%

probability Return probability*(Return-Expected Return)^2
0.1 -50 0.1*(-50-10.9)^2=370.881
0.2 -6 0.2*(-6-10.9)^2=57.122
0.4 9 0.4*(9-10.9)^2=1.444
0.2 30 0.2*(30-10.9)^2=72.962
0.1 75 0.1*(75-10.9)^2=410.881
Total=913.29%

Standard deviation=[Total probability*(Return-Expected Return)^2/Total probability]^(1/2)

=(913.29)^(1/2)

=30.22%(Approx).